OpenStudy (anonymous):

a linear equation when the slope is -.5 and f(-2)=4

OpenStudy (anonymous):

find the equation of the line with \(m=.5\) through the point \((-2,4)\)

OpenStudy (anonymous):

start with \[y-y_1=m(x-x_1)\] putting \(m=.5,x_1=-2,y_1=4\)